At what temperature will water freeze on a standard, single car driveway with 0.50 cm thick ice (assume 105 L of water) that has been salted with 11.0 kg CaCl2? *AȚf K.cm i, K,--1.86 Cim, assume d = 1 g/mL
Answer #1

Water will freeze at a temperature of -5.27oC

Explanation

Volume of water = 105 L

Mass of water = (volume of water) * (density of water)

Mass of water = (105 L) * (1 kg/L)

Mass of water = 105 kg

mass CaCl2 = 11.0 kg = 11000 g

moles CaCl2 = (mass CaCl2) / (molar mass CaCl2)

moles CaCl2 = (11000 g) / (111 g/mol)

moles CaCl2 = 99.1 mol

molality CaCl2 = moles CaCl2 / mass of water (kg)

molality CaCl2 = 99.1 mol / 105 kg

molality CaCl2 = 0.944 m

According to equation for depression in freezing point,

\small \DeltaTf = = i * Kf * m

where \small \Delta Tf = decrease in freezing point

i = Van't Hoff factor = number of ions after dissociation = 3 (for CaCl2)

Kf = -1.86 oC/m

m = molality of solute = 0.944 m

Substituting the values

\small \DeltaTf = (3) * (-1.86 oC/m) * (0.944 m)

\small \DeltaTf = -5.27 oC

Water will freeze at a temperature of -5.27oC